# A Calculus of Overlays

### v1

@@ Appendix C Multi-Path Self-Reference in Scala
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2—but DOT’s self variable is single-valued, making this intersection unrealizable.
@@
This is well-defined in Overlay-Calculus. Inside $\mathrm{MyInner}$, the reference $\mathrm{MyOuter}.\mathbf{this}$ has de Bruijn index $n=1$ (one scope level up from $\mathrm{MyInner}$ to $\mathrm{MyOuter}$). When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths—one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$—and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
Appendix D Multi-Path Self-Reference in the NixOS Module System
This appendix demonstrates that the NixOS module system rejects the multi-path self-reference pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits (unevaluated module values imported into other fixpoints) and submoduleWith for objects (evaluated in their own fixpoints).
@@ Appendix D Multi-Path Self-Reference in the NixOS Module System
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error—it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error (Appendix C): both systems assume that each option (or type member) has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In Overlay-Calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix E Well-Definedness of the Semantic Functions
@@ E.3. Monotonicity
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}$. The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
@@
$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|(\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
@@
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\;(n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
@@
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension—i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
E.4. Existence of the Least Fixed Point
Theorem E.2.
@@ E.4. Existence of the Least Fixed Point
By Proposition E.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined—the analogue of divergence. ∎
Appendix F Böhm Tree Correspondence: Proofs
@@ F.1. Böhm Trees
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\text{if }M\to^{*}_{h}\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Hyland and Ong, 2000).
F.2. Path Encoding
@@
F.3. Single-Path Lemma
@@
Lemma F.1 (Single Path).
@@
This is a reference, not a record. The $\mathtt{this}$ resolution walks up $n$ scope levels. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
@@ F.4. Substitution Lemma
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{this}$ resolves the reference by walking up $n$ scope levels to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma F.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
@@
Then $\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}$. The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
@@ F.5. Convergence Preservation
Lemma F.3 (Result-step).
@@
The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma F.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma F.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
@@
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma F.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$. ∎
Theorem F.4 (Convergence preservation).
@@
A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus). In both cases, Lemma F.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
F.6. Adequacy
@@
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. By Lemma F.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
@@
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$. In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value). By Lemma F.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
F.7. Full Abstraction

### v2

@@ Appendix B Multi-Path Self-Reference in Other Systems
The rejection is not a surface-level restriction but a consequence of DOT’s type-theoretic foundations (Amin et al., 2016). In DOT, an object is constructed as $\nu(x:T)\,d$, where the self variable $x$ binds to a single object. Object1.MyInner and Object2.MyInner are distinct path-dependent types, each constraining the outer self-reference to a different object. Their intersection requires outer to simultaneously return Object1 and Object2, but DOT’s self variable is single-valued, making this intersection unrealizable.
@@
This is well-defined in overlay-calculus. The reference $\mathrm{MyOuter}.\mathbf{this}$ inside $\mathrm{outer}$ has de Bruijn index $n=1$: starting from $\mathrm{outer}$’s enclosing scope $\mathrm{MyInner}$, one $\mathtt{this}$ step reaches $\mathrm{MyOuter}$. When $\mathrm{MyObjectA}$ inherits from both $\mathrm{Object1}.\mathrm{MyInner}$ and $\mathrm{Object2}.\mathrm{MyInner}$, the $\mathtt{this}$ function (equation 6) resolves $\mathrm{MyOuter}.\mathbf{this}$ by searching through $\mathtt{supers}(\mathrm{MyObjectA})$ for overlay paths matching $\mathrm{MyInner}$’s definition site. It finds two inheritance-site paths, one through $\mathrm{Object1}$ and one through $\mathrm{Object2}$, and returns both. Both paths lead to records that inherit from the same $\mathrm{MyOuter}$, so querying $\mathrm{MyObjectA}.\mathrm{outer}$ yields the $\mathtt{properties}$ of $\mathrm{MyOuter}$ regardless of which route is taken. Since inheritance is idempotent, the two routes contribute the same properties and no disambiguation is needed.
NixOS Module System
The NixOS module system rejects the same pattern with a static error. The translation uses the module system’s own abstractions: deferredModule for traits, which are unevaluated module values imported into other fixpoints, and submoduleWith for objects, which are evaluated in their own fixpoints.
@@
The rejection occurs because Object1.MyInner and Object2.MyInner each carry the full staticModules of the MyInner deferred module, including the declaration options.outer. When MyObjectA imports both, the module system’s mergeOptionDecls encounters two declarations of the same option and raises a static error; it cannot express that both declarations originate from the same trait and should be unified.
The rejection is analogous to Scala’s “conflicting base types” error above: both systems assume that each option or type member has a single declaration site, and reject the multi-path situation where two inheritance routes introduce the same declaration independently. In overlay-calculus, the same pattern is well-defined because $\mathtt{overlays}$ (equation 3) recognizes that both routes lead to the same definition, and $\mathtt{supers}$ (equation 2) collects both inheritance-site contexts without duplication.
Appendix C Well-Definedness of the Semantic Functions
@@ C.3. Monotonicity
$\mathtt{properties}(p)=\bigl\{\,\ell\bigm|(\_,p_{\mathrm{overlay}})\in\mathtt{supers}(p),\;\ell\in\mathtt{defines}(p_{\mathrm{overlay}})\,\bigr\}.$
The function $\mathtt{supers}$ appears as the set being iterated over ($\in\mathtt{supers}(p)$). The predicate $\ell\in\mathtt{defines}(\cdot)$ involves only the AST primitive $\mathtt{defines}$, not any mutually defined function. Enlarging $\mathtt{supers}(p)$ can only add pairs, hence can only add labels to the result. Positive in $\mathtt{supers}$.
@@
$\mathtt{overlays}(p)=\{p\}\cup\bigl\{\,p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)\bigm|{}\\ (\_,p_{\mathrm{branch}})\in\mathtt{supers}(\mathtt{init}(p)),\;\mathtt{last}(p)\in\mathtt{defines}(p_{\mathrm{branch}})\,\bigr\}.$
@@
$\mathtt{bases}(p)=\bigl\{\,p_{\mathrm{target}}\bigm|p_{\mathrm{overlay}}\in\mathtt{overlays}(p),\\ (n,\ell_{\mathrm{down},*})\in\mathtt{inherits}(p_{\mathrm{overlay}}),\;p_{\mathrm{target}}\in\mathtt{resolve}(\ldots)\,\bigr\}.$
@@
In summary, every mutually defined function appears on the right-hand side of every equation only inside the pattern $x\in f(\ldots)$ within a set comprehension, i.e., as the domain of an existential quantifier. No complement, set difference, or negated membership test appears in any equation. Therefore $F$ is a monotone operator on the complete lattice $D$. ∎
C.4. Existence of the Least Fixed Point
Theorem C.2.
@@ C.4. Existence of the Least Fixed Point
By Proposition C.1, $F$ is monotone on the complete lattice $D$. By the Knaster–Tarski theorem (Tarski, 1955), $F$ has a least fixed point $\vec{d}^{*}=\bigsqcup_{n\geq 0}F^{n}(\bot)$, where $\bot$ assigns the empty set to every input. The components of $\vec{d}^{*}$ are the semantic functions of Section 3.
Concretely, the least fixed point is the union of the finite approximation chain: $\bot\sqsubseteq F(\bot)\sqsubseteq F^{2}(\bot)\sqsubseteq\cdots$. A membership $\ell\in\mathtt{properties}(p)$ holds in $\vec{d}^{*}$ if and only if it appears at some finite stage $F^{k}(\bot)$, which corresponds to a finite derivation tree using the equations. When no finite derivation exists, the membership is absent from the least fixed point and the query is undefined, the analogue of divergence. ∎
Appendix D Böhm Tree Correspondence: Proofs
@@ D.1. Böhm Trees
$\mathrm{BT}(M)=\begin{cases}\bot&\text{if }M\text{ has no HNF}\\[6.0pt] \lambda x_{1}\ldots x_{m}.\;y\bigl(\mathrm{BT}(M_{1}),\;\ldots,\;\mathrm{BT}(M_{k})\bigr)&\begin{aligned} &\text{if }M\to^{*}_{h}\\[-2.0pt] &\lambda x_{1}\ldots x_{m}.\;y\;M_{1}\cdots M_{k}\end{aligned}\end{cases}$
Böhm tree equivalence $\mathrm{BT}(M)=\mathrm{BT}(N)$ is the standard observational equivalence of the lazy $\lambda$-calculus (Barendregt, 1984; Abramsky and Ong, 1993).
D.2. Path Encoding
@@
D.3. Single-Path Lemma
@@
Lemma D.1 (Single Path).
@@
This is a reference, not a record. The $\mathtt{this}$ function walks up $n$ steps from the enclosing scope. By the inductive hypothesis on the enclosing translations, each step of $\mathtt{this}$ encounters exactly one matching pair in $\mathtt{supers}$.
@@ D.4. Substitution Lemma
Then $\mathcal{T}(x)=\uparrow^{\!n}.\mathrm{argument}$ (where $n$ is the de Bruijn index of $x$), and $C\mathbin{\triangleright}\mathrm{result}$ contains this reference. In the inherited tree $C$, $\mathtt{resolve}$ and $\mathtt{this}$ resolve the reference by navigating from the reference’s enclosing scope up $n$ steps to the binding $\lambda$, where the inheritance $\{\mathcal{T}(\lambda x.\,x),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$ overlays the $\mathrm{argument}$ slot with $\mathcal{T}(V)$. Since $C$ is the $\mathrm{tail}$ subtree of $\mathcal{T}\bigl((\lambda x.\,M)\;V\bigr)$, which is a closed ANF term, Lemma D.1 applies and $\mathtt{this}$ finds exactly one path, so $\mathtt{resolve}$ returns $\mathcal{T}(V)$’s subtree. Since $M[V/x]=V$, we have $\mathcal{T}(M[V/x])=\mathcal{T}(V)$, and the properties coincide.
@@
Then
$\mathcal{T}(M)=\{z\mapsto\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\},\;\mathrm{result}\mapsto\mathcal{T}(M^{\prime})\}.$
The substitution distributes: $M[V/x]=\mathbf{let}\;z=V_{1}[V/x]\;V_{2}[V/x]\;\mathbf{in}\;M^{\prime}[V/x]$. By induction on $V_{1}$, $V_{2}$, and $M^{\prime}$, the properties under each subtree ($z$ and $\mathrm{result}$) coincide, since references to $x$ inside each subtree resolve to $\mathcal{T}(V)$ via the same $\mathtt{this}$ mechanism.
@@ D.5. Convergence Preservation
Lemma D.3 (Result-step).
@@
The $\mathrm{result}$ label at the root is defined as $\mathrm{tail}.\mathrm{result}$, so $\text{root}\mathbin{\triangleright}\mathrm{result}$ resolves to the $\mathrm{result}$ path inside the inheritance $\mathrm{tail}=\{\mathcal{T}(\lambda x.\,M),\;\mathrm{argument}\mapsto\mathcal{T}(V)\}$. This inheritance is exactly the inherited tree $C$ of Lemma D.2, and $\mathrm{tail}\mathbin{\triangleright}\mathrm{result}$ is $C\mathbin{\triangleright}\mathrm{result}$. By Lemma D.2, the properties under $C\mathbin{\triangleright}\mathrm{result}$ coincide with those of $\mathcal{T}(M[V/x])$.
@@
Since $\mathrm{result}\mapsto\mathcal{T}(M^{\prime})$ is a property definition, the subtree at $\text{root}\mathbin{\triangleright}\mathrm{result}$ is $\mathcal{T}(M^{\prime})$. Inside $\mathcal{T}(M^{\prime})$, each reference to $x$ is $x.\mathrm{result}$, which resolves to the $\mathrm{result}$ path inside the $x$ subtree $\{\mathcal{T}(\lambda y.\,M^{\prime\prime}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$. This is the inherited tree $C$ of Lemma D.2, so $x.\mathrm{result}$ has the same properties as $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. In the $\lambda$-calculus, $M=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$, and the head reduct is $N=M^{\prime}[(V_{1}\;V_{2})/x]$. In $\mathcal{T}(N)$, each occurrence of $x$ is replaced by $(V_{1}\;V_{2})$, whose result, by the tail-call variant above, also yields $\mathcal{T}(M^{\prime\prime}[V_{2}/y])$. Therefore the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$ for all $p$. ∎
Theorem D.4 (Convergence preservation).
@@
A head reduction step in ANF takes one of two forms: a tail call $(\lambda x.\,M^{\prime})\;V\to_{h}M^{\prime}[V/x]$, or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}\to_{h}M^{\prime}[(V_{1}\;V_{2})/x]$ (since $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}=(\lambda x.\,M^{\prime})(V_{1}\;V_{2})$ in the $\lambda$-calculus). In both cases, Lemma D.3 (tail-call and let-binding variants respectively) shows that the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n}$ in $\mathcal{T}(M)$ coincide with the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}^{n-1}$ in $\mathcal{T}(N)$ for all $n\geq 1$, where $N$ is the head reduct. Hence the abstraction shape appears at depth $n$ in the pre-reduct if and only if it appears at depth n1n-1 in the post-reduct. ∎
D.6. Adequacy
@@
Inductive step ($k\geq 1$): $M$ is not an abstraction, so $M$ is either a tail call or a $\mathbf{let}$-binding, both of which are $\beta$-redexes. Let $N$ be the head reduct, so $M\to_{h}N\to^{k-1}_{h}\lambda x.\,M^{\prime}$. By Lemma D.3 (tail-call or let-binding variant), the properties at $\text{root}\mathbin{\triangleright}\mathrm{result}\mathbin{\triangleright}p$ in $\mathcal{T}(M)$ coincide with those at $\text{root}\mathbin{\triangleright}p$ in $\mathcal{T}(N)$. By the induction hypothesis, $\mathcal{T}(N)$ converges at depth $\leq k-1$, so $\mathcal{T}(M)$ converges at depth $\leq k$.
@@
Inductive step ($n\geq 1$): the abstraction shape appears at depth $n$ but not at depth 0. By the base case argument, $M$ is not an abstraction, so $M$ is either a tail call $V_{1}\;V_{2}$ or a $\mathbf{let}$-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M^{\prime}$. In both cases, $M$ is a $\beta$-redex (since $V_{1}=\lambda x.\,M^{\prime}$ or $V_{1}=\lambda y.\,M^{\prime\prime}$ respectively, the only closed ANF value). By Lemma D.3 (tail-call or let-binding variant respectively), the properties at depth n1n-1 in $\mathcal{T}(N)$ match those at depth $n$ in $\mathcal{T}(M)$, where $N$ is the head reduct. So $\mathcal{T}(N)$ converges at depth n1n-1. By the induction hypothesis, $N$ has a head normal form, and since $M\to_{h}N$, so does $M$. ∎
D.7. Full Abstraction